OpenStudy (anonymous):

how do you solve 3/x+3 = 5/2x+6 + 1/x-2

OpenStudy (lalaly):

\[(\frac{3}{(x+3)} \times \frac{2}{2}) - \frac{5}{2x+6}=\frac{1}{x-2}\]\[\frac{6}{2x+6}-\frac{5}{2x+6}=\frac{1}{x-2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

==> x=-8

OpenStudy (lalaly):

\[\frac{1}{2x+6}=\frac{1}{x-2}\]\[2x+6=x-2\]\[2x-x=-2-6\]\[2x=-8\]\[x=-4\]